Given the line 2x – 3y = 9 and the point (4, –1), find lines through the point that are

(a) parallel to the given line and
(b) perpendicular to it.

Found 2 solutions by josgarithmetic, stanbon:
Answer by josgarithmetic(39616)

Your given line is some standard form, 2x-3y=c.
c in the given equation is a constant, but it can be used as an unknown variable.

Line parallel is of the form, 2x-3y=c. Find c using the point given to be contained in the line.

Line perpendicular is of the form 3x+2y=c. Find c using the point given to be contained in the line.

Answer by stanbon(75887)

Given the line 2x – 3y = 9 and the point (4, –1), find lines through the point that are
-------------------
Find the slope of the give line:
3y = 2x-9
y = (2/3)x - 3
slope = 2/3
--------------------------------
(a) parallel to the given line
slope must be (2/3)
--
Form:: y = mx + b
Solve for "b" using the point and slope.
-1 = (2/3)4 + b
-3/3 = 8/3 + b
b = -11/3
Equation:
y = (2/3)x - (11/3)
=============================
(b) perpendicular to it
slope must be -3/2
Form: y = mx + b
Solve for "b" using the point and slope
-1 = (-3/2)4 + b
-1 = -6 = b
b = 5
Equation:
y = (-3/2)x + 5
-----------------------
Cheers,
Stan H.
====================